OpenStudy (anonymous):

Employed women. If 60% of all women are employed outside of home, find the probability that in a sample of 20 women, A. Exactly 15 are employed B. At least 10 are employed C. At most 5 are not employed outside the home

OpenStudy (ybarrap):

You can use the binomial distribution A. $$ p=0.6\\ q=1-p\\ {20\choose15}p^{15}q^{5} $$ B. $$ \sum _{n=10}^{20}{20\choose n}p^{n}q^{20-n} $$ C B. $$ \sum _{n=6}^{20}{20\choose n}p^{n}q^{20-n} $$ Make sense?